# Two datasets are better than one: Method of double moments for 3-D reconstruction in cryo-EM

### v1

@@ Abstract
Cryo-electron microscopy (cryo-EM) is a powerful imaging technique for reconstructing three-dimensional molecular structures from noisy tomographic projection images of randomly oriented particles. We introduce a new data fusion framework, termed the method of double moments (MoDM), which reconstructs molecular structures from two instances of the second-order moment of projection images obtained under distinct orientation distributions—one uniform, the other non-uniform and unknown. We prove that these moments generically uniquely determine the underlying structure, up to a global rotation and reflection, and we develop a convex-relaxation-based algorithm that achieves accurate recovery using only second-order statistics. Our results demonstrate the advantage of collecting and modeling multiple datasets under different experimental conditions, illustrating that leveraging dataset diversity can substantially enhance reconstruction quality in computational imaging tasks.
@@ 1 Introduction
Cryo-electron microscopy has become an increasingly popular technique for single-particle three-dimensional structure reconstruction [23, 35, 3, 14]. It aims to recover the underlying molecular structure from a large collection of noisy 2-D tomographic projection images taken at unknown and typically random orientations, with applications in structural biology, medicine, and drug discovery [49, 62, 20, 54, 67].
A fundamental challenge in cryo-EM is the high noise level and the unknown random viewing directions of particles in the sample, which complicate the reconstruction process. This paper introduces a new statistical approach for 3-D reconstruction that relies solely on low-order statistics—specifically, two instances of the second-order moment of the observed projection images, each computed from images with different distributions of orientations. Crucially, we prove that under mild conditions, these moments uniquely determine the underlying structure (up to a global rotation and reflection), and we devise a computationally efficient algorithm based on convex relaxation to perform the reconstruction.
@@
An important practical motivation for our method comes from the fact that projection orientations in cryo-EM are typically not uniformly distributed. This anisotropy in orientation distribution can result from experimental factors such as preferred particle orientations or sample preparation artifacts [61, 41, 24, 47, 15, 43, 4]. To address this issue, recent experimental advances have sought to manipulate particle orientation distributions to be more uniform—for example, by tilting the specimen [2], applying laser flash melting [60], adding charged detergents [40], introducing ultrasonic excitation [65], encapsulating molecules in liposomes [72] or in protein shells (“nanocrates”) [32], or by high-speed droplet vitrification before protein diffusion [28].
Our proposed method, termed the method of double moments (MoDM), is naturally suited to such experimental scenarios. It leverages two datasets collected under distinct orientation distributions—one ideally uniform, made possible through experimental interventions, and the other non-uniform and unknown a priori. Although this setting is increasingly relevant in practice, it has not been systematically explored in existing statistical reconstruction frameworks. Our results demonstrate that fusing such complementary datasets enables exact structure recovery using only second-order statistics, thereby opening new avenues for robust and efficient cryo-EM reconstruction despite variations in particle orientations.
@@ 1.1 Notation
Denote by $\mathbb{Z}_{\geq 0}$ the set of nonnegative integers, and by $\mathbf{1}_{\mathcal{X}}$ the indicator function of a set $\mathcal{X}$. For a matrix $M$, denote by $M^{\mathsf{T}}$ and $M^{\mathsf{H}}$ its transpose and conjugate transpose, respectively. We denote by $\mathsf{O}(n)$ the set of real orthogonal matrices of dimension $n$, that is, $\mathsf{O}(n)=\{M\in\mathbb{R}^{n\times n}:MM^{\mathsf{T}}=M^{\mathsf{T}}M=I\}$. Similarly, $\mathsf{SO}(n)$ denotes the special orthogonal group of dimension $n$, consisting of real orthogonal matrices with determinant +1+1, i.e., $\mathsf{SO}(n)=\{M\in\mathsf{O}(n):\det(M)=1\}$. We also use $\mathsf{O}(n,\mathbb{C})$ to denote the set of complex orthogonal matrices of dimension $n$, defined by $\mathsf{O}(n,\mathbb{C})=\{M\in\mathbb{C}^{n\times n}:MM^{\mathsf{T}}=M^{\mathsf{T}}M=I\}$. Note that, unlike unitary matrices, the orthogonality condition here does not involve complex conjugation.
@@
For a function $f:\mathbb{R}^{d}\to\mathbb{R}$, we use the following convention for the Fourier transform:
$\widehat{f}(\boldsymbol{\omega})=\int_{\mathbb{R}^{d}}f(\mathbf{x})\,e^{-2\pi i\,\mathbf{x}\cdot\boldsymbol{\omega}}\,d\mathbf{x},$
@@ 1.2 Image formation model
In the proposed algorithm, the projection images $\{{I_{i}}\}_{i=1}^{N}$ are used solely to estimate the second-order moment of the underlying structure. The influence of the point spread function can be compensated during this estimation, provided that the functions $h_{i}$, for $i=1,\ldots,N$, exhibit sufficiently non-overlapping zero crossings [55, 44]. Similarly, in-plane shifts can be corrected using established centering algorithms [29]. We therefore assume that these corrections have been performed in advance and omit the point spread functions and shifts in the subsequent analysis.
@@ 1.3 Existing methods
Since this is a classical example of an incomplete data problem, the expectation-maximization (EM) algorithm is commonly employed to maximize the likelihood [17]. Popular state-of-the-art methods resort to EM-based procedures—commonly referred to as 3-D iterative refinement—which alternate between estimating the posterior distribution over rotations (E-step) and updating the 3-D structure estimate by maximizing the expected log-likelihood (M-step) [56, 51, 48].
@@
The method of moments offers several advantages over competing approaches, such as maximum likelihood-based methods with 3-D iterative refinement. One key benefit is that empirical moments in (1.7) can be computed using only one or two passes over the data, after which the raw dataset no longer needs to be accessed [71, 9, 55, 44]. This property significantly reduces computational costs for sufficiently large datasets, especially when compared to the repeated access and high iteration count required in likelihood-based refinement pipelines. Another advantage lies in its applicability to small molecular structures (e.g., below 40 kDa), where state-of-the-art software implementations of 3-D iterative refinement encounter challenges [51, 48]. Even in these challenging regimes, particle locations can still be detected reliably from micrographs [63, Figure 10f–h], which allows empirical moments to be formed and the method of moments to be applied—bypassing the limitations of traditional refinement.
Historically, the method of moments was introduced to the cryo-EM setting by Kam [33], who observed numerically that third-order population moments can uniquely determine bandlimited molecular structures (to be precisely defined in Section 2.1)—under the assumption that the viewing directions are uniformly distributed over $\mathsf{SO}(3)$. These observations have since been rigorously justified under various technical assumptions [5, 22, 21]. Despite these strengths, significant challenges remain. Accurate estimation of the $d$th-order moment requires a number of samples that scales as $\omega(\sigma^{2d})$ [5, 1], which becomes prohibitive even at moderate moment orders such as $d=3$, particularly in high-noise settings. Furthermore, when moments are discretized as tensors, both storage and computational costs grow exponentially with the moment order, presenting serious practical limitations. Recent approaches attempt to mitigate these issues by exploiting compressed low-rank tensor formats [30].
@@ 1.4 Contributions
  • We prove that the molecular structure is uniquely identified by the two second-order moments (rather than up to a finite ambiguity), provided that the structure is bandlimited (see Section 2.4 for a precise statement). This result significantly improves prior identifiability guarantees that rely on higher-order moments or strong generative assumptions.
@@ 2.1 Basis representation of structure
Here we assume that $\widehat{\Phi}$ is bandlimited, in the sense that $\widehat{\Phi}$ can be represented exactly in the form in (2.1) for a finite value of $L$. Although this is an idealization, should this not hold, higher components $A_{\ell}^{m}(r)Y_{\ell}^{m}(\theta,\Phi)$ for $\ell>L$ can be treated as an additional source of approximation error. An illustration of this bandlimiting assumption and its effect on the reconstructed structure is provided in Figure 2.1.
Section (2.3) presents explicit analytical expressions for the moment $m_{2}[\Phi,\rho]$ for bandlimited functions of the form (2.1), using the complex-valued spherical harmonics introduced above. When describing our algorithm, it will however be convenient to also use real-valued spherical harmonics. We therefore introduce notation for converting complex-valued spherical harmonics basis coefficients into coefficients in the real-valued spherical harmonics basis.
@@ 2.2 Basis representation of rotation distribution
We model the distribution of viewing orientations as having a density function $\rho(R)$ with respect to the uniform/Haar measure over the rotation group $\mathsf{SO}(3)$. To facilitate analysis and to reflect symmetries commonly present—or enforceable via data augmentation—in cryo-EM experiments, we work under the following two structural assumptions on $\rho(R)$:
@@
The complex-valued basis coefficients $B_{p,u}$ satisfy constraints from the real-valued nature and normalization of $\rho$:
@@ 3 Algorithm: Method of Double Moments
These steps are detailed in Sections 3.13.3, respectively. Algorithm 1 presents the complete procedure in summary form. Here, all quantities with a widetilde $\widetilde{\cdot}\,$ are empirically accessible from the observed samples.
@@
Compute $\widetilde{C}$ in (3.2) through numerical quadrature, and obtain $\widetilde{A}_{\ell}$ from $\widetilde{C}_{\ell}$ via Cholesky decomposition followed by taking the real part.
@@ 3.1 Step 1: Kam’s method via uniform sample moment
Kam [33] showed, via the Fourier slice theorem, that if the viewing directions are uniformly distributed over $\mathsf{SO}(3)$, then the autocorrelation function of the 3-D volume with itself can be recovered directly from the second-order moment of 2-D projection images. In particular, for a uniform distribution $\rho_{1}^{*}$, it follows from (2.18) and the orthogonality of the Legendre polynomials that, for each $\ell=0,\ldots,L$ and $r,r^{\prime}\in[0,r_{\text{max}}]$,
@@
Assume we have access to samples on the grid in (2.19). We approximate the integral in (3.1) by first using the sample moment $\widetilde{m}_{2}[\Phi^{*},\rho_{1}^{*}]$ instead of the population moment, and second by using numerical quadrature with quadrature points $\varphi_{s},\varphi_{q}$ and quadrature weights $w_{s,q}$, for $s,q=0,\ldots,M_{\varphi}$, i.e., we compute
@@
respectively, where $\{r_{1},\ldots,r_{M_{r}}\}$ denote the sampling radial points. Inserting these definitions into (3.1) and (3.2) gives
@@
Recall from Remark 2.1 that $\widecheck{A}_{\ell}$ is real for even $\ell$ and purely imaginary for odd $\ell$. Motivated by this structure, applying the Cholesky decomposition of the matrix $C_{\ell}$, we obtain factors
@@
In practice, we replace $C_{\ell}$ with its sample counterpart $\widetilde{C}_{\ell}$, perform a Cholesky decomposition, and take the real part to obtain an estimate $\widetilde{A}_{\ell}$ for $\breve{A}_{\ell}$. The remaining task is to recover the orthogonal matrices $\{O_{\ell}\}$, which would in turn allow us to fully estimate $\{A_{\ell}\}$.
@@ 3.2 Step 2: Formulate optimization with the non-uniform sample moment
where we denote by $\breve{A}_{\ell}(r)$ the row of $\breve{A}_{\ell}$ corresponding to the radial sample $r$.
@@
This naturally leads to a least-squares problem, where one seeks to find the matrices $\{O_{\ell}\}$ and parameters $\{B_{p,u}\}$ by minimizing
@@ 3.3 Step 3: Convex relaxation and alternating refinement
By relaxing the orthogonality constraint on $\mathcal{O}$, we introduce an unconstrained surrogate
$\mathcal{X}=\text{blockdiag}_{\ell=0,\ldots,L}(X_{\ell})$
for $\mathcal{O}$, where each $X_{\ell}$ has the same matrix size as $O_{\ell}$. We then solve for $\mathcal{X}$ in a least-squares sense and subsequently orthogonalize the result to recover $\mathcal{O}$. Specifically, we restrict the search to the following set:
@@
where the equality follows from the block-diagonal structures of $\mathcal{X}$ and $\mathcal{O}$. Each term in the sum can be solved efficiently using the orthogonal Procrustes procedure, which admits a closed-form solution by computing the singular value decomposition (SVD) of each diagonal block $X_{\ell}$ [52].
@@ 3.5 Numerical experiments
For varying $M$, we generate $120000$ projection images of size $M\times M$, with signal-to-noise-ratio $0.1$ and affected by convolution with point spread functions with $10000$ defocus groups. The projection images were generated with voltage 300 $kV$, defocus values ranging between 11 $\mu m$ and 33 $\mu m$ and spherical aberration $2$ mm. The ground truth distribution of viewing angles was taken to be a mixture of $8$ von Mises-Fisher-Langevin distributions.
The result of running Algorithm 1 on bandlimited structures is shown in Figures 3.2 and 3.2. Notably, the algorithm successfully recovers bandlimited functions up to a maximum bandlimit $L(M)$ that depends on the number of pixels in the projection images. This stems from the inversion of the linear system in Eq. (3.11), which is only well-posed for a limited range of bandlimits $L$, for a fixed value of $M$.

### v2

@@ Abstract
Cryo-electron microscopy (cryo-EM) is a powerful imaging technique for reconstructing three-dimensional molecular structures from noisy tomographic projection images of randomly oriented particles. We introduce a new data fusion framework, termed the method of double moments (MoDM), which reconstructs molecular structures from two instances of the second-order moment of projection images obtained under distinct orientation distributions: one uniform, the other non-uniform and unknown. We prove that these moments generically uniquely determine the underlying structure, up to a global rotation and reflection, and we develop a convex-relaxation-based algorithm that achieves accurate recovery using only second-order statistics. Our results demonstrate the advantage of collecting and modeling multiple datasets under different experimental conditions, illustrating that leveraging dataset diversity can substantially enhance reconstruction quality in computational imaging tasks.
@@ 1 Introduction
Cryo-electron microscopy has become an increasingly popular technique for single-particle three-dimensional (3-D) structure reconstruction [23, 35, 3, 14]. It aims to recover the underlying molecular structure from a large collection of noisy two-dimensional (2-D) tomographic projection images taken at unknown and typically random orientations, with applications in structural biology, medicine, and drug discovery [49, 62, 20, 54, 67].
A fundamental challenge in cryo-EM is the high noise level and the unknown random viewing directions of particles in the sample, which complicate the reconstruction process. This paper introduces a new statistical approach for 3-D reconstruction that relies solely on low-order statistics, specifically, two instances of the second-order moment of the observed projection images, each computed from images with different distributions of orientations. Crucially, we prove that under mild conditions, these moments uniquely determine the underlying structure (up to a global rotation and reflection), and we devise a computationally efficient algorithm based on convex relaxation to perform the reconstruction.
@@
An important practical motivation for our method comes from the fact that projection orientations in cryo-EM are typically not uniformly distributed. This anisotropy in orientation distribution can result from experimental factors such as preferred particle orientations or sample preparation artifacts [61, 41, 24, 47, 15, 43, 4]. To address this issue, recent experimental advances have sought to manipulate particle orientation distributions to be more uniform, by, for example, tilting the specimen [2], applying laser flash melting [60], adding charged detergents [40], introducing ultrasonic excitation [65], encapsulating molecules in liposomes [72] or in protein shells (“nanocrates”) [32], or by high-speed droplet vitrification before protein diffusion [28].
Our proposed method, termed the method of double moments (MoDM), is naturally suited to such experimental scenarios. It leverages two datasets collected under distinct orientation distributions: one ideally uniform, made possible through experimental interventions, and the other non-uniform and unknown a priori. Although this setting is increasingly relevant in practice, it has not been systematically explored in existing statistical reconstruction frameworks. Our results demonstrate that fusing such complementary datasets enables exact structure recovery using only second-order statistics, thereby opening new avenues for robust and efficient cryo-EM reconstruction despite variations in particle orientations.
@@ 1.1 Notation
Denote by $\mathbb{R}$ the set of real numbers, by $\mathbb{Z}_{\geq 0}$ the set of nonnegative integers, and by $\mathbf{1}_{\mathcal{X}}$ the indicator function of a set $\mathcal{X}$. For a matrix $M$, denote by $M^{\mathsf{T}}$ and $M^{\mathsf{H}}$ its transpose and conjugate transpose, respectively. We denote by $\mathsf{O}(n)$ the set of real orthogonal matrices of dimension $n$, that is, $\mathsf{O}(n)=\{M\in\mathbb{R}^{n\times n}:MM^{\mathsf{T}}=M^{\mathsf{T}}M=I\}$. Similarly, $\mathsf{SO}(n)$ denotes the special orthogonal group of dimension $n$, consisting of real orthogonal matrices with determinant +1+1, i.e., $\mathsf{SO}(n)=\{M\in\mathsf{O}(n):\det(M)=1\}$. We also use $\mathsf{O}(n,\mathbb{C})$ to denote the set of complex orthogonal matrices of dimension $n$, defined by $\mathsf{O}(n,\mathbb{C})=\{M\in\mathbb{C}^{n\times n}:MM^{\mathsf{T}}=M^{\mathsf{T}}M=I\}$. Note that, unlike unitary matrices, the orthogonality condition here does not involve complex conjugation.
@@
For a square-integrable function $f:\mathbb{R}^{d}\to\mathbb{R}$, we use the following convention for the Fourier transform:
$\widehat{f}(\boldsymbol{\omega})=\int_{\mathbb{R}^{d}}f(\mathbf{x})\,e^{-2\pi i\,\mathbf{x}\cdot\boldsymbol{\omega}}\,d\mathbf{x},\qquad\boldsymbol{\omega}\in\mathbb{R}^{d},$
@@ 1.2 Image formation model
In the proposed algorithm, the projection images $\{{I_{i}}\}_{i=1}^{N}$ are used solely to estimate the second-order moment of the underlying structure. The influence of the point spread function can be compensated for during this estimation, provided that the functions $h_{i}$, for $i=1,\ldots,N$, exhibit sufficiently non-overlapping zero crossings [55, 44]. Similarly, in-plane shifts can be corrected using established centering algorithms [29]. We therefore assume that these corrections have been performed in advance and omit the point spread functions and shifts in the subsequent analysis.
@@ 1.3 Existing methods
Since this is a classical example of an incomplete data problem, the expectation-maximization (EM) algorithm is commonly employed to maximize the likelihood [17]. Popular state-of-the-art methods resort to EM-based procedures (commonly referred to as 3-D iterative refinement) which alternate between estimating the posterior distribution over rotations (E-step) and updating the 3-D structure estimate by maximizing the expected log-likelihood (M-step) [56, 51, 48].
@@
The method of moments offers several advantages over competing approaches, such as maximum likelihood-based methods with 3-D iterative refinement. One key benefit is that empirical moments in (1.7) can be computed using only one or two passes over the data, after which the raw dataset no longer needs to be accessed [71, 9, 55, 44]. This property significantly reduces computational costs for sufficiently large datasets, especially when compared to the repeated access and high iteration count required in likelihood-based refinement pipelines. Another advantage lies in its applicability to small molecular structures (e.g., below 40 kDa), where state-of-the-art software implementations of 3-D iterative refinement encounter challenges [51, 48]. Even in these challenging regimes, particle locations can still be detected reliably from micrographs [63, Figure 10f–h], which allows empirical moments to be formed and the method of moments to be applied, bypassing the limitations of traditional refinement.
Historically, the method of moments was introduced to the cryo-EM setting by Kam [33], who observed numerically that third-order population moments can uniquely determine bandlimited molecular structures (to be precisely defined in Section 2.1), under the assumption that the viewing directions are uniformly distributed over $\mathsf{SO}(3)$. These observations have since been rigorously justified under various technical assumptions [5, 22, 21]. Despite these strengths, significant challenges remain. Accurate estimation of the $d$th-order moment requires a number of samples that scales as $\omega(\sigma^{2d})$ [5, 1], which becomes prohibitive even at moderate moment orders such as $d=3$, particularly in high-noise settings. Furthermore, when moments are discretized as tensors, both storage and computational costs grow exponentially with the moment order, presenting serious practical limitations. Recent approaches attempt to mitigate these issues by exploiting compressed low-rank tensor formats [30].
@@ 1.4 Contributions
  • We prove that the molecular structure is uniquely identified by the two second-order moments and their corresponding first-order moments (rather than up to a finite ambiguity), provided that the structure is bandlimited. See Section 2.4 for a precise statement. This result significantly improves prior identifiability guarantees that rely on higher-order moments or strong generative assumptions.
@@ 2.1 Basis representation of structure
Here we assume that $\widehat{\Phi}$ is bandlimited, in the sense that $\widehat{\Phi}$ can be represented exactly in the form in (2.1) for a finite value of $L$. Although this is an idealization, should this not hold, higher components $A_{\ell}^{m}(r)Y_{\ell}^{m}(\theta,\Phi)$ for $\ell>L$ can be treated as an additional source of approximation error. An illustration of this bandlimiting assumption is provided in Figure 2.1.
Section 2.3 presents explicit analytical expressions for the moment $m_{2}[\Phi,\rho]$ for bandlimited functions of the form (2.1), using the complex-valued spherical harmonics introduced above. When describing our algorithm, it will however be convenient to also use real-valued spherical harmonics. We therefore introduce notation for converting complex-valued spherical harmonics basis coefficients into coefficients in the real-valued spherical harmonics basis.
@@ 2.2 Basis representation of rotation distribution
We model the distribution of viewing orientations as having a density function $\rho(R)$ with respect to the uniform/Haar measure over the rotation group $\mathsf{SO}(3)$. To facilitate analysis and to reflect symmetries commonly present (or enforceable via data augmentation) in cryo-EM experiments, we work under the following two structural assumptions on $\rho(R)$:
@@
The complex-valued basis coefficients $B_{p,u}$ satisfy constraints from the real-valuedness and normalization of $\rho$:
@@ 3 Algorithm: Method of Double Moments
These steps are detailed in Sections 3.13.3, respectively. Algorithm 1 presents the complete procedure in summary form. Here, all quantities with a tilde $\widetilde{\cdot}\,$ are empirically accessible from the observed samples.
@@
Compute $\widetilde{C}$ in (3.2) through numerical quadrature, and obtain $\widetilde{A}_{\ell}$ from $\widetilde{C}_{\ell}$ via Cholesky decomposition.
@@ 3.1 Step 1: Kam’s method via uniform sample moment
Kam [33] showed, via the Fourier slice theorem, that if the viewing directions are uniformly distributed over $\mathsf{SO}(3)$, then the autocorrelation function of the 3-D volume with itself can be computed directly from the second-order moment of 2-D projection images. In particular, for a uniform distribution $\rho_{1}^{*}$, it follows from (2.18) and the orthogonality of the Legendre polynomials that, for each $\ell=0,\ldots,L$ and $r,r^{\prime}\in[0,r_{\text{max}}]$,
@@
Assume we have access to samples on the grid in (2.19). We approximate the integral in (3.1) by firstly using the sample moment $\widetilde{m}_{2}[\Phi^{*},\rho_{1}^{*}]$ instead of the population moment, and secondly by using numerical quadrature with quadrature points $\varphi_{s},\varphi_{q}$ and quadrature weights $w_{s,q}$, for $s,q=0,\ldots,M_{\varphi}$, i.e., we compute
@@
respectively, where $\{r_{1},\ldots,r_{M_{r}}\}$ denote the radial sampling points. Inserting these definitions into (3.1) and (3.2) gives
@@
Recall from Remark 2.1 that $\widecheck{A}_{\ell}$ is real for even $\ell$ and purely imaginary for odd $\ell$. Applying the Cholesky decomposition of the matrix $C_{\ell}$, we then obtain factors
@@
In practice, we replace $C_{\ell}$ with its sample counterpart $\widetilde{C}_{\ell}$, and perform a Cholesky decomposition, to obtain an estimate $\widetilde{A}_{\ell}$ for $\breve{A}_{\ell}$. The remaining task is to recover the orthogonal matrices $\{O_{\ell}\}$, which would in turn allow us to fully estimate $\{A_{\ell}\}$.
@@ 3.2 Step 2: Formulate optimization with the non-uniform sample moment
where we denote by $\breve{A}_{\ell}(r)$ the row of $\breve{A}_{\ell}$ corresponding to the radius $r$.
@@
This naturally leads to a least-squares problem, where we seek to find the matrices $\{O_{\ell}\}$ and parameters $\{B_{p,u}\}$ by minimizing
@@ 3.3 Step 3: Convex relaxation and alternating refinement
By relaxing the orthogonality constraint on $\mathcal{O}$, we introduce a relaxed variable
$\mathcal{X}=\text{blockdiag}_{\ell=0,\ldots,L}(X_{\ell}),$
where each $X_{\ell}$ has the same matrix size as $O_{\ell}$. We then solve for $\mathcal{X}$ in a least-squares sense and subsequently orthogonalize the result to recover $\mathcal{O}$. Specifically, we restrict the search to the following set:
@@
where the equality follows from the block-diagonal structures of $\mathcal{X}$ and $\mathcal{O}$. Each term in the sum can be minimized efficiently using the orthogonal Procrustes procedure, which admits a closed-form solution by computing the singular value decomposition (SVD) of each diagonal block $X_{\ell}$ [52].
@@ 3.5 Numerical experiments
For varying $M$, we generate $100,000$ projection images of size $M\times M$, with signal-to-noise-ratio $0.1$ and affected by convolution with point spread functions with $20,000$ defocus groups. The projection images were generated with voltage 300 $kV$, defocus values ranging between 11 $\mu m$ and 33 $\mu m$, amplitude contrast $0.1$ and spherical aberration $2$ mm. The ground truth distribution of viewing angles was taken to be a mixture of $8$ von Mises-Fisher-Langevin distributions.
A representative result of running Algorithm 1 on bandlimited structures is shown in Figures 3.2 and 3.2. Notably, we typically observe that the algorithm successfully recovers bandlimited functions up to a maximum bandlimit $L(M)$ that depends on the number of pixels in the projection images, provided a sufficient number of projection images are used to reduce variance when estimating the moments and the distribution of viewing angles is sufficiently non-uniform (making this quantitatively precise is left to future work). The figure shows the maximum bandlimits $L(M)$ with observed successful reconstruction for image sizes $M\in\{64,128,256\}$ together with the unsuccessful reconstruction for $L(M)+1$. However, lower bandlimits can result in unsuccessful reconstruction for some viewing distributions. This stems from the inversion of the linear system in Eq. (3.11), which is only well-posed for a limited range of bandlimits $L$, for a fixed value of $M$, and the fact that properties of the optimization landscape depends on the viewing angle distribution coefficients $\{B_{p,u}\}$. In particular, we intuitively expect the optimization landscape to require a distribution of viewing angles that is sufficiently distinct from the uniform distribution, but not so distinct that it e.g. has regions of near-zero probability.